At a distance of 100 m and parallel to a river, a side channel flows through...

  1. At a distance of 100 m and parallel to a river, a side channel flows through the river’s floodplain. The bed of the river and side channel are at the same elevation. At present, the depth of water in the river is 3 m and the depth of water in the side channel is 1 m. Assume the floodplain is composed of silty sand with a hydraulic conductivity (K) of 10-5 m/s with no confining layers. Compute the discharge per unit width (Q*) of water flowing through the floodplain in m2/day.
Answer #1

sofi - Here Hydraullic gradient, i= A4 or i= 3-1 = 0.09 Also k=los mis 0.02 we know - Know - Kit Adext S = 10*x0.02 x lou 2 -